There were 16 fans at Martin's soccer game. The home team had 3 times as many fans as the visiting team. How many fans did the home team have?
1
Expert's answer
2020-01-31T12:30:46-0500

X+3*X=16

4*X=16

X=4

3*X = 12


Answer: the home team has 12 fans.